# Discovery and Analysis of Afterglows from Poorly Localised GRBs with the Gravitational-wave Optical Transient Observer (GOTO) All-sky Survey

### v1

@@ Acknowledgements
51GEPI, Observatoire de Paris, Universit´e PSL, CNRS, 5 Place Jules Janssen, Meudon, 92190, France
52Institute of Physics and Technology, Institutskiy Pereulok, 9, Dolgoprudny 141701, Russia.
53Astronomical Institute Anton Pannekoek, University of Amsterdam, PO Box 94249, NL-1090 GE Amsterdam, the Netherlands
54Materials Science and Applied Mathematics, Malmö University, SE-205 06 Malmö, Sweden
@@
55Nordic Optical Telescope, Rambla José Ana Fernández Pérez 7, ES-38711 Breña Baja, Spain
@@
56E. Kharadze Georgian National Astrophysical Observatory, Mt. Kanobili, Abastumani 0301, Adigeni, Georgia.
57Department of Physics, George Washington University, 725 21st St NW, Washington, DC 20052, USA.
58LUX, Observatoire de Paris, Université PSL, CNRS, Sorbonne Université, Meudon, 92190, France

### v2

@@ Acknowledgements
52GEPI, Observatoire de Paris, Universite´ PSL, CNRS, 5 Place Jules Janssen, F-92190 Meudon, France
53Institute of Physics and Technology, Institutskiy Pereulok, 9, Dolgoprudny 141701, Russia
54Astronomical Institute Anton Pannekoek, University of Amsterdam, PO Box 94249, NL-1090 GE Amsterdam, the Netherlands
55Materials Science and Applied Mathematics, Malmö University, SE-205 06 Malmö, Sweden
@@
56Nordic Optical Telescope, Rambla José Ana Fernández Pérez 7, ES-38711 Breña Baja, Spain
@@
57E. Kharadze Georgian National Astrophysical Observatory, Mt. Kanobili, Abastumani 0301, Adigeni, Georgia
58Department of Physics, George Washington University, 725 21st St NW, Washington, DC 20052, USA
59LUX, Observatoire de Paris, Université PSL, CNRS, Sorbonne Université, Meudon, 92190, France